# Uncertainty-Calibrated Test-Time Model Adaptation without Forgetting

### v1

@@ 6 Conclusion
  • In Section B, we show more experimental results to compare the out-of-distribution performance, calibration and efficiency with state-of-the-art methods on ImageNet-C with different corruption types.
  • In Section C, we give more experimental results to demonstrate the anti-forgetting ability of our EATA.
  • In Section D, we provide more discussions on related training-time robustification studies.
@@ A.2 More Details on Datasets
Following the settings of Tent [12] and MEMO [14], we conduct experiments on four benchmark datasets for out-of-distribution generalization, i.e., CIFAR-10-C, ImageNet-C [9], ImageNet-R [65], and ACDC [66].
CIFAR-10-C and ImageNet-C consist of corrupted versions of the validation images on CIFAR-10 [72] and ImageNet [73], respectively. The corruptions include 15 diverse types of 4 main categories (i.e., noise, blur, weather, and digital). Each corruption type has 5 different levels of severity.
@@
ImageNet-R contains 30,000 images with various artistic renditions of 200 ImageNet classes, which are primarily collected from Flickr and filtered by Amazon MTurk annotators.
@@
ACDC contains four categories of images collected in adverse conditions, including fog, night, rain, and snow. Following [64], we use 400 unlabeled images from each adverse condition for adaptation.
A.3 More Experimental Protocols
Our EATA and EATA-C.
@@ A.3 More Experimental Protocols
Following Tent [12] and CoTTA [64], we use ResNet-26 for CIFAR-10 experiments, ResNet-50 and ViTBase for ImageNet experiments, and Segformer-B5 for ACDC experiments. In classification experiments, the models are trained on the original CIFAR-10 or ImageNet training set and then tested on clean or the aforementioned OOD test sets. In semantic segmentation experiments, the model is trained on Cityscapes and then tested on ACDC. For a fair comparison, the parameters of ViTBase and Segformer-B5 are directly obtained from the timm [^1] https://github.com/pprp/timm and Segformer [68], respectively. ResNet-50 is trained via the official code in the torchvision [^2] https://github.com/pytorch/vision library with stochastic depth. ResNet-26 is trained via the official code of TTT by the same hyper-parameters, replacing the group norm with the batch norm, and removing the rotation head. For test time adaptation, we use SGD as the update rule, with a momentum of 0.9 and a batch size of 64. In EATA and ETA, the learning rate is set to 0.005 on CIFAR-10, 0.00025/0.001 for ResNet-50/ViTBase on ImageNet, and 7.5×$\times 10^{-5}$ on ACDC, respectively (following Tent, SAR and CoTTA). The entropy constant $E_{0}$ in Eqn. (3) is set to $0.4\times\ln C$ in classification experiments and $0.1\times\ln C$ in segmentation, where $C$ is the number of task classes. The $\epsilon$ in Eqn. (8) is set to 0.4 for CIFAR-10, and 0.05 for ImageNet and ACDC. The trade-off parameter $\beta$ in Eqn. (10) is set to 1/2,000/500 for CIFAR-10/ImageNet/ACDC to make two losses have a similar magnitude. In EATA-C, the learning rate is set to 0.005/0.1 for ResNet-50/ViTBase on ImageNet, and 0.0005 on ACDC, respectively. The sub-network is obtained via stochastic depth regularization [63] with a drop ratio of 0.2/0.6 for ImageNet/ACDC. The entropy constant $E_{0}$ in Eqn. (3) is set to 0.4/0.5×$\times\ln C$, and the $\epsilon$ in Eqn. (8) is set to 0.05/0.07 for ViTBase/ResNet-50 on ImageNet. In Eqn. (16), the trade-off parameter $\alpha$ is set to 50/500 on ImageNet/ACDC and $\beta$ is set to 0.1 on ImageNet. Note that for the semantic segmentation task, our EATA-C does not perform min-max entropy regularization and active sample selection due to the long-tailed problem [74, 75]. For both EATA and EATA-C, the moving average factor $\alpha$ in Eqn. (6) is set to 0.1, and we use 2,000/20 samples for ImageNet/ACDC to calculate $\omega(\theta_{i})$ in Eqn. (11).
@@
Compared Methods. For TTA [76], BN adaptation [32], MEMO [14], CoTTA [64], and SAR [18] the hyper-parameters follow their original papers or MEMO. Specifically, the augmentation size of TTA [76] is set to 32 and 64 for CIFAR-10 and ImageNet, respectively. For BN adaptation [32], both the batch size $B$ and prior strength $N$ are set to 64. The learning rate in CoTTA [64] for ViTBase is set to 0.005, and the augmentation threshold $p_{th}$ is set to 0.1. Other hyper-parameter settings of CoTTA, and the hyper-parameter settings of MEMO [14] and SAR [18] can be found in their original paper. For Tent [12], we use SGD as the update rule with a momentum of 0.9. The batch size is 64 for both ImageNet and CIFAR-10 experiments. The learning rate is set to 0.00025/0.001 for ResNet-50/ViTBase on ImageNet and 0.005 on CIFAR-10, respectively. Note that the hyper-parameters of Tent are totally the same as our EATA for a fair comparison. For TTT [11], we strictly follow their original settings except for the augmentation size at test time for ImageNet experiments. According to TTT’s implementation, the augmentation size is set to 64, which, however, is very time-consuming (e.g., about 12 GPU hours on a single Tesla V100 GPU on ImageNet-C with a specific corruption type and severity level). In our implementation, we decrease this augmentation size to 20, which has only a slight performance difference compared with augmentation size 64. For example, the performances of TTT on ImageNet-C (Gaussian noise, severity level 5) with ResNet-18 are 26.2% (64) vs. 26.0% (20). We recommend the original papers of the above methods to readers for more implementation details.
Appendix B More Results on Out-of-distribution Performance, Efficiency and Calibration
@@
In Table D, we provide more results to compare our EATA and EATA-C with state-of-the-art methods in the single-domain adaptation scenario on ImageNet-C with the severity level 5. Although the compared methods exhibit stronger performance in the less challenging adaptation scenario, our EATA and EATA-C still demonstrate substantial superiority. Specifically, by filtering the unreliable and redundant samples from adaptation, our EATA achieves significant accuracy improvement while requiring less backward-propagation (e.g., 55.5% (72,446) vs.61.8% (32,524) over the best method SAR on ViTBase). Meanwhile, EATA-C further enhances performance while mitigating calibration error on both ResNet-50 and ViTBase, verifying the promising effectiveness of our EATA and EATA-C across different adaptation scenarios.
In Table E, we provide the complete results for the lifelong test-time adaptation on ACDC. Compared with Tent, our EATA effectively alleviates error accumulation by excluding unreliable gradients and constraining important weights from drastic changes. Furthermore, besides maintaining strong performance in the long term, our EATA-C is able to further accumulate knowledge to improve performance over adaptation, demonstrating strong potential.
In Figure B, we show the number of backward propagation of our ETA on ImageNet-C with different corruption types and severity levels. Across various corruption types, our ETA shows great superiority over existing methods in terms of adaptation efficiency. Compared with MEMO (50,000$\times$64) and Tent (50,000), our ETA only requires 31,741 backward passes (averaged over 15 corruption types) when the severity level is set to 3. The reason is that we exclude some unreliable and redundant test samples out of test-time optimization. In this case, we only need to perform backward computation on those remaining test samples, leading to improved efficiency.
In Table F, we compared ETA and EATA with existing methods on CIFAR-10-C. Under the same base model (ResNet-26 with batch normalization), ETA achieves a lower average error than Tent (19.4% vs. 20.2%) with fewer requirements of back-propagation (8,192 vs. 10,000). Moreover, the performance gain over the base model of ETA is larger than that of TTT, i.e., +9.0% vs. +7.2% on accuracy. These results are consistent with the ones on ImageNet-C/R, further demonstrating the effectiveness and superiority of our method.
Comparison with Tent using Different Learning Rates. In our sample-adaptive weight $S{({\bf x})}$ in Eqn. (8), each test sample has a specific weight $S{({\bf x})}$ and the value of $S{({\bf x})}$ is always larger than 1. In this sense, training with sample-adaptive weight $S{({\bf x})}$ indeed has the same effect as training with larger learning rates. Therefore, we compare our EATA with the baseline (Tent) using different learning rates. We increase the learning rate from $2.5\times 10^{-4}$ (which is the default of Tent) to $25.0\times 10^{-4}$ and report results in Table A.
With the learning rate increasing from $2.5\times 10^{-4}$ to $10.0\times 10^{-4}$, the error of Tent decreases from 45.3% to 43.9%, indicating that a larger learning rate may enhance the performance in some cases. However, when the learning rate becomes larger to $20.0\times 10^{-4}$, the performance of Tent degrades. More critically, our EATA method outperforms Tent with varying learning rates. These results verify that simply enlarging the learning rate is not able to achieve competitive performance with our proposed sample-adaptive adaptation method, demonstrating our superiority.
TABLE A: Comparison with Tent under different learning rates (×$\times 10^{-4}$) on ImageNet-C (Gaussian noise) regarding Error (%).
EATA (ours) Tent [12]
Severity $lr=2.5$ $lr=2.5$ $lr=10.0$ $lr=20.0$ $lr=25.0$
Level 3 42.6 45.3 43.9 44.4 45.1
Level 5 65.0 71.6 72.2 83.6 87.1
EATA under Different Random Orders. Table B records EATA’s performance (mean & stdev.) on randomly shuffled test samples with 10 different random seeds (from 2020 to 2029). From the results, EATA performs consistently across different random orders, showing the stability of EATA.
TABLE B: The mean and stdev. of corruption accuracy (%) of EATA over 10 random orders, on ImageNet-C (level 5) with ResNet-50.
Gauss. Shot Impul. Defoc. Glass Motion Zoom Avg.
34.9±${}_{\pm 0.2}$ 36.9±${}_{\pm 0.1}$ 35.8±${}_{\pm 0.2}$ 33.6±${}_{\pm 0.3}$ 33.3±${}_{\pm 0.2}$ 47.2±${}_{\pm 0.3}$ 52.7±${}_{\pm 0.1}$ 35.8±${}_{\pm 0.2}$
EATA with Different Numbers of Test Samples. We investigate the effect of the total number of test samples ($N$) in EATA, where fewer samples are sampled from the entire test set. From Table C, EATA works well and consistently outperforms Tent counterpart, regardless of the number of test samples. Meanwhile, if there are many test samples, EATA would benefit more, i.e., larger accuracy gain.
TABLE C: Comparison with Tent [12] w.r.t. corruption accuracy (%) with fewer ($N$) test samples on ImageNet-C (Gaussian noise, level 5) with ResNet-50.
Method $N\small{=}$256 $N\small{=}$512 $N\small{=}$1,024 $N\small{=}$2,048 $N\small{=}$4,096 $N\small{=}$10,000
Tent 13.7 18.2 14.8 15.2 17.5 20.9
EATA 14.8 20.1 16.3 18.7 23.5 27.6
Appendix C More Results on Prevent Forgetting
@@
In this section, we provide more results to demonstrate the effectiveness of our EATA in preventing forgetting. We report the comparison results of EATA (lifelong) vs. Tent (lifelong) and EATA vs. Tent in Figures C and D, respectively. In the lifelong adaptation scenario, Tent suffers more severe ID performance degradation than that of reset adaptation (i.e., Figure D), showing that the more optimization steps, the more severe forgetting. Moreover, with the increase of the severity level, the ID clean accuracy degradation of Tent increases accordingly. This result indicates that the OOD adaptation with more severe distribution shifts will result in more severe forgetting. In contrast, our methods achieve higher OOD corruption accuracy and, meanwhile maintain the ID clean accuracy (competitive to the original accuracy that is tested before any OOD adaptation) in both two adaptation scenarios (reset and lifelong). These results are consistent with that in the main paper and further demonstrate the effectiveness of our proposed anti-forgetting weight regularization.
Appendix D More Discussions on Related Training-Time Robustification
To defend against distribution shifts, many prior studies seek to enlarge the training data distribution to enable it to cover the possible shift that might be encountered at test time, such as adversarial training strategies [77, 78, 79], various data augmentation techniques [80, 81, 82, 65, 83] and searching/enhancing sub-networks of a deep model [84, 85]. However, it is hard to anticipate all possible test shifts at training time. In contrast, we seek to conquer this test distribution shift by directly learning from test data.
@@ Appendix D More Discussions on Related Training-Time Robustification
TABLE D: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, normal-$\uparrow$) and Expected Calibration Error(%,normal-$\downarrow$). “BN” and “LN” denote batch and layer normalization. All results are evaluated in the single-domain adaptation scenario (i.e., the model parameters are reset before adapting to a new corruption type) except for MEMO [14]. We use * to denote episodic adaptation.
@@
Model Method Metric Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brit. Contr. Elastic Pixel JPEG Avg. # Forwards # Backwards
@@
EATA (Ours) Acc. 35.6 37.8 36.1 35.3 35.4 47.6 53.1 51.4 45.9 60.1 67.6 45.8 58.2 60.3 55.4 48.36
ECE 10.5 10.2 10.3 12.9 12.3 8.3 6.6 6.7 8.2 4.7 4.5 9.9 5.1 5.3 5.6 8.08 50,000 26,188
EATA-C (Ours) Acc. 37.2 39.2 38.0 35.5 35.9 48.2 52.3 51.9 46.5 60.4 67.3 48.7 58.0 60.5 56.0 49.04
ECE 7.1 6.8 6.8 9.3 8.9 4.3 3.6 3.0 4.5 1.8 1.9 4.5 2.3 2.4 2.4 4.63 82,492 32,492
@@
EATA (Ours) Acc. 50.5 51.3 51.3 55.9 56.0 60.7 58.0 64.2 62.2 70.1 77.4 66.7 63.9 70.5 68.9 61.84
ECE 10.6 10.1 10.3 9.4 8.8 7.8 9.1 6.4 7.3 5.4 2.8 6.1 6.6 4.5 4.8 7.34 50,000 32,524
EATA-C (Ours) Acc. 56.6 57.6 57.2 57.8 59.8 64.6 64.4 69.1 67.2 73.3 78.6 67.0 70.2 73.9 71.2 65.90
ECE 5.3 4.8 5.2 5.3 4.1 3.5 4.0 2.8 2.7 2.1 1.4 2.9 2.2 1.7 2.2 3.34 82,364 32,364
TABLE E: Semantic segmentation results (mIoU in %) on the Cityscapes-to-ACDC in lifelong test-time adaptation scenario. The model is continually adapted to the four adverse conditions for ten rounds without resetting model parameters, i.e., lifelong adaptation. All results are evaluated based on the Segformer-B5 model.
@@
TABLE F: Comparison on CIFAR-10-C. Results are evaluated in the single-domain adaptation scenario, and averaged over 15 different corruption types and 5 severity levels (totally 75). We use * to denote episodic adaptation.
Model Acc. (%) #Forwards #Backwards
ResNet-26 (GroupNorm) 77.5 10,000 0
[…]
ResNet-26 (GroupNorm)+JT 77.2 10,000 0
[…]
ResNet-26 (BatchNorm) 71.6 10,000 0
@@
Figure B: Comparison between ETA and Tent in terms of the number of backward propagation on ImageNet-C with different corruption types and severity levels.
Figure C: Comparison of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50. We record the OOD corruption accuracy on each corrupted test set and the associated ID clean accuracy (after OOD adaptation). The model performs lifelong adaptation, in which the model parameters will never be reset.
Figure D: Comparisons of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50. We record the OOD corruption accuracy on each corrupted test set and the associated ID clean accuracy (after OOD adaptation). The model parameters of both Tent and our EATA are reset before adapting to a new corruption type.

### v2

@@ 6 Conclusion
Acknowledgments
This work was partially supported by the Joint Funds of the National Natural Science Foundation of China (Grant No.U24A20327), and TCL Science and Technology Innovation Fund, China.
@@ Acknowledgments
  • In Section B, we present additional experimental results and ablation studies to further demonstrate our superiority in out-of-distribution performance, calibration, efficiency, and stability.
  • In Section C, we provide further insights, and empirical/visual evidence to validate the design choices in EATA-C.
@@ A.2 More Details on Datasets
Following the settings of Tent [12] and MEMO [14], we conduct experiments on three benchmark datasets for out-of-distribution generalization, i.e., ImageNet-C [9], ImageNet-R [66], and ACDC [67].
ImageNet-C consists of various versions of corruption applied to 50,000 validation images from ImageNet. The dataset encompasses 15 distinct corruption types of 4 main categories, including Gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, fog, brightness, contrast, elastic transformation, pixelation, and JPEG compression. Each corruption type is characterized by 5 different levels of severity, with higher severity levels indicating a more severe distribution shift.
@@
ImageNet-R contains 30,000 images with various artistic renditions of 200 ImageNet classes. These images are primarily collected from Flickr and filtered by Amazon MTurk annotators.
@@
ACDC contains four categories of images collected in adverse conditions, including fog, night, rain, and snow. Following CoTTA [63], we use 400 unlabeled images from each adverse condition for continuous test-time adaptation.
A.3 More Experimental Protocols on Evaluation
@@
Following Tent [12] and CoTTA [63], we use ResNet-50 and ViT-Base for ImageNet experiments, and Segformer-B5 for ACDC experiments. In classification experiments, the models are trained on the original ImageNet training set and then tested on clean or the aforementioned OOD test sets. In semantic segmentation experiments, the model is trained on Cityscapes and then tested on ACDC. For a fair comparison, the parameters of ViTBase and Segformer-B5 are directly obtained from the timm [^1] https://github.com/pprp/timm and Segformer [^2] https://github.com/NVlabs/SegFormer [68], respectively. ResNet-50 is trained via the official code in the torchvision [^3] https://github.com/pytorch/vision library with stochastic depth.
[…]
@@
Compared Methods. For BN adaptation [29], MEMO [14], CoTTA [63], SAR [18], ROID [64], Rdump [71], TEA [65], and DAT [70], the hyper-parameters follow their original papers or MEMO. Specifically, for BN adaptation [29], both the batch size $B$ and prior strength $N$ are set to 64. The learning rate in CoTTA [63] for ViTBase is set to 0.005, and the augmentation threshold $p_{th}$ is set to 0.1. TEA [65] uses SGD with a learning rate of 0.00025/0.005 for ResNet-50/ViT-Base, respectively. Other hyper-parameter settings of CoTTA and TEA, as well as those for MEMO, SAR, ROID, Rdump, and DAT, can be found in their original paper. For Tent [12], we use SGD with a momentum of 0.9. The batch size is 64 for both ImageNet experiments and the learning rate is set to 0.00025/0.001 for ResNet-50/ViTBase, respectively. Note that the hyper-parameters of Tent are the same as our EATA for a fair comparison.
Appendix B More Experimental Results
B.1 Results under Single-Domain Adaptation
@@
In Table A, we provide additional results to evaluate the effectiveness of our EATA and EATA-C under single-domain adaptation scenarios. From the results, while the compared methods exhibit stronger performance in the less challenging adaptation setting, our EATA and EATA-C still demonstrate substantial superiority. Specifically, by filtering the unreliable and redundant samples from adaptation, our EATA achieves substantial accuracy improvement while requiring significantly fewer backpropagations, e.g., on ViT-Base, average accuracy increases from 55.5% (SAR) to 61.8% (EATA) with a reduction in backward passes from 72,446 to 32,524. Meanwhile, EATA-C further enhances performance and reduces calibration error on both ResNet-50 and ViTBase. These results are consistent with our findings in lifelong adaptation (Table II), verifying the effectiveness of our EATA and EATA-C across diverse adaptation scenarios.
B.2 More Results on Semantic Segmentation
In Table B, we present the complete results for lifelong test-time adaptation on the ACDC dataset. Although lifelong adaptation offers exposure to a broader range of data with the potential for improved performance, Tent and DAT tend to suffer from error accumulation and catastrophic forgetting, leading to a decline in subsequent adaptations. In contrast, our EATA mitigates these issues by excluding unreliable gradients and constraining important weights from drastic changes, thereby maintaining a more stable performance. Moreover, our EATA-C is able to accumulate learned knowledge from prior adaptation and showcase consistent improvement over lifelong adaptation, increasing the average mIOU on four adverse datasets from 59.8% (first round) to 62.3% (tenth round). These results further highlight the long-term effectiveness of our methods in leveraging the potential of lifelong test-time adaptation.
B.3 Results under Diverse Severity Levels
[…]
B.4 More Results on Prevent Forgetting
@@
We provide more results to demonstrate the effectiveness of our EATA in preventing forgetting. We report the comparison results of EATA (lifelong) vs. Tent (lifelong) and EATA vs. Tent in Figures F and G, respectively. In the lifelong adaptation scenario, Tent suffers more severe ID performance degradation than that of reset adaptation (i.e., Figure G), showing that the more optimization steps, the more severe forgetting. Moreover, with the increase of the severity level, the ID clean accuracy degradation of Tent increases accordingly. This result indicates that the OOD adaptation with more severe distribution shifts will result in more severe forgetting. In contrast, our methods achieve higher OOD corruption accuracy and, meanwhile maintain the ID clean accuracy (competitive to the original accuracy that is tested before any OOD adaptation) in both two adaptation scenarios (reset and lifelong). These results are consistent with those in the main paper and further demonstrate the effectiveness of our proposed anti-forgetting weight regularization.
B.5 More Ablation Studies on EATA-C’s Components
EATA-C aims to achieve a favorable balance among accuracy, calibration, and efficiency. We supplement more ablation studies on different datasets and architectures to verify the effectiveness of each component, as in Table C. Our findings on ImageNet-R with ResNet-50 are consistent with those in Table VI, including 1) Using the consistency loss alone already yields a strong baseline; 2) Adding the entropy regularizer and Fisher regularizer further improves TTA accuracy while reducing ECE; 3) By filtering out unreliable and redundant samples, active sample selection significantly boosts computational efficiency while maintaining or improving accuracy.
Moreover, we provide further explanations regarding the performance variations between ViT-Base and ResNet-50 when incorporating active sample selection. This is primarily attributed to the quality of pseudo labels for TTA. ViT-Base generally yields higher-quality pseudo labels, with consistently higher accuracy and lower ECE. Considering adaptation efficiency, the active sample selection strategy may remove a few instances that could have been beneficial to adaptation, resulting in a small drop in accuracy (e.g., from 66.4% to 65.9%), while reducing the required backward passes by over 35%. In contrast, ResNet-50 tends to produce less stable pseudo labels, which could adversely affect the performance if all samples are used indiscriminately, as evidenced by the collapse of TEA [65] on ResNet in Table II. In this case, active sample selection not only reduces unnecessary backward passes, but also enhances accuracy by focusing test-time adaptation on more trustworthy samples.
Appendix C Additional Discussions
C.1 Foundations of Disagreement-Based Data Uncertainty Indicator
In EATA-C, we devise a disagreement-based indicator to identify uncertain samples. This is inspired by margin-based theories in machine learning [61, 62], which indicate that samples near decision boundaries are inherently more uncertain and have been well justified with theoretical guarantees. To be specific, we leverage prediction disagreement between the full network and sub-network to detect these uncertain boundary samples, where conflicting predictions are more likely to occur. Figure C provides a toy illustration of this principle and we depict our model and data uncertainty based on Figure C below. 1) Model uncertainty. In our method, activating only a subset of parameters results in a new decision boundary for the sub-model, e.g., Sub-Model 1/2. Our consistency loss aligns the sub-model boundary with that of the full model, thereby aiming to minimize the model uncertainty and obtain an optimal decision boundary. 2) Data uncertainty. Unlike our consistency loss that measures model uncertainty, in which samples across the entire data space may yield low consistency loss, data uncertainty is reflected more prominently through the disagreement of predictions near the decision boundary.
Moreover, we provide empirical evidence of using prediction disagreement as the data uncertainty indicator in Table VIII and Figure C. From the results, we reveal that a substantial number of samples (over 13,000) exhibit persistent prediction disagreement even after extensive adaptation. Notably, about 90% of these samples are misclassified by the well-trained model, suggesting intrinsic noise or ambiguity near decision boundaries, aligned with the definition of data uncertainty. Collectively, the margin-based theories and our empirical results provide a sound basis for adopting disagreement as a practical indicator of data uncertainty in TTA.
Figure B: Toy illustration. Uncertain data near the decision boundary tend to cause disagreed predictions among the sub- and full model.
Figure C: Evolution of uncertain samples (i.e., with disagreed predictions) during adaptation on ImageNet-C (Gauss, level 5) using Eqn. (10).
C.2 Parameter Update Strategies: Norm-Only vs. Full Updates
In our main experiments, EATA-C updates only the normalization layers by following established practices of existing TTA work [12, 65]. Here, we further validate this design choice by comparing it against a variant that updates all model parameters, namely EATA-C . The results in Table D highlight the following. 1) Reduced Network Divergence: Updating the normalization layer effectively decreases the divergence loss between the full network and the sub-network, from 0.565 (source model) to 0.451; 2) Improved Accuracy and Calibration: While updating all model parameters yields a marginally lower divergence, our EATA-C that updates only the norm layers leads to higher accuracy (56.8% vs. 56.0%) and lower ECE (5.2% vs. 5.9%). This improvement is mainly attributed to the increased stability of TTA, where restricting updates to the normalization layers acts as a regularizer for online adaptation and reduces the risk of forgetting and overfitting by keeping the majority of parameters frozen, also known as parameter-efficient tuning. 3) Lower Overhead: Our EATA-C also demonstrates reduced adaptation time (114.9s vs. 140.9s) and memory usage (5786.6MB vs. 9920.5MB) compared to EATA-C, making it more practical for real-world applications.
C.3 Visualization of Reliable and Unreliable Test Samples
We show some selected and removed test samples for TTA according to our entropy-based indicator in Eqn. (3). From Figure E, the selected samples generally exhibit higher image quality (i.e., with clear and distinguishable class features), where the model can generate more reliable pseudo labels for adaptation. In contrast, the removed samples are often ambiguous even for human interpretation, leading to potentially incorrect predictions that would introduce noisy learning signals when used for adaptation. This highlights the effectiveness of our entropy-based identification in improving the robustness of TTA.
@@ C.3 Visualization of Reliable and Unreliable Test Samples
TABLE A: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, $\uparrow$) and Expected Calibration Error($\downarrow$). “BN” and “LN” denote batch and layer normalization. All results are evaluated in the single-domain adaptation scenario (i.e., the model parameters are reset before adapting to a new corruption type) except for MEMO [14]. We use * to denote episodic adaptation.
@@
Acc. 35.6 37.8 36.1 35.3 35.4 47.6 53.1 51.4 45.9 60.1 67.6 45.8 58.2 60.3 55.4 48.36
EATA (Ours) ECE 10.5 10.2 10.3 12.9 12.3 8.3 6.6 6.7 8.2 4.7 4.5 9.9 5.1 5.3 5.6 8.08 50,000 26,188
Acc. 37.2 39.2 38.0 35.5 35.9 48.2 52.3 51.9 46.5 60.4 67.3 48.7 58.0 60.5 56.0 49.04
EATA-C (Ours) ECE 7.1 6.8 6.8 9.3 8.9 4.3 3.6 3.0 4.5 1.8 1.9 4.5 2.3 2.4 2.4 4.63 82,492 32,492
@@
Acc. 50.5 51.3 51.3 55.9 56.0 60.7 58.0 64.2 62.2 70.1 77.4 66.7 63.9 70.5 68.9 61.84
EATA (Ours) ECE 10.6 10.1 10.3 9.4 8.8 7.8 9.1 6.4 7.3 5.4 2.8 6.1 6.6 4.5 4.8 7.34 50,000 32,524
Acc. 56.6 57.6 57.2 57.8 59.8 64.6 64.4 69.1 67.2 73.3 78.6 67.0 70.2 73.9 71.2 65.90
EATA-C (Ours) ECE 5.3 4.8 5.2 5.3 4.1 3.5 4.0 2.8 2.7 2.1 1.4 2.9 2.2 1.7 2.2 3.34 82,364 32,364
TABLE B: Semantic segmentation results (mIoU in %) on the Cityscapes-to-ACDC in lifelong TTA scenario. The model is continually adapted to the four adverse conditions for ten rounds without resetting model parameters, i.e., lifelong adaptation. All results are evaluated based on the Segformer-B5 model.
@@
TABLE C: Effects of components in EATA-C. Results obtained in single-domain TTA scenario, i.e., the model parameters are reset before adapting to a new corruption. CL denotes consistency loss. ER denotes entropy regularization. FR denotes Fisher regularization. $S$ denotes active sample selection.
Experiment CL ER FR $S$ ResNet-50 over ImageNet-R
Acc. ECE #Forwards #Backwards
Source 38.0 17.7 30,000 0
@@
[…]
7 46.9 14.4 35,147 5,147
8 44.6 11.9 34,991 4,991
9 (ETA-C) 46.8 13.4 35,137 5,137
10 47.1 13.7 35,179 5,179
11 44.6 12.1 34,935 4,935
12 (EATA-C) 47.1 13.3 35,122 5,122
Experiment CL ER FR $S$ ViT-Base over ImageNet-C (level 5)
Acc. ECE #Forwards #Backwards
Source 39.8 7.5 50,000 0
[…]
7 65.6 4.4 83,182 33,182
8 52.9 3.4 73,323 23,323
9 (ETA-C) 65.3 3.4 82,015 32,015
10 66.2 4.3 83,429 33,429
11 53.2 3.4 73,723 23,723
12 (EATA-C) 65.9 3.3 82,364 32,364
TABLE D: Comparison of parameter update strategies. EATA-C (Ours) updates only the normalization layer parameters, while EATA-C updates all model parameters. Results are obtained on ImageNet-C (Gaussian, level 5) with ViT-Base. Loss denotes the average KL-Divergence between the full network and the sub-network across the dataset during online TTA. Time and peak memory usage are measured by processing 50,000 images.
Method
Loss
Acc. (%)
ECE (%)
Time (s)
Memory (MB)
Source 0.565 12.9 14.2 55.7 816.6
SAR - 43.1 8.0 167.3 7433.2
EATA-C 0.437 56.0 5.9 140.9 9920.5
EATA-C (Ours) 0.451 56.8 5.2 114.9 5786.6
Figure D: Comparison between ETA and Tent w.r.t. the number of backward passes on ImageNet-C with different corruption types and severity levels.
Figure E: Visualization of selected (reliable) and removed (unreliable) test samples for adaptation according to Eqn. (3). Predicition confidence is the highest class probability. Results are obtained on ImageNet-C(Gaussian, level 5) using the source ViT-Base model.
Figure F: Comparison of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50. We record the OOD corruption accuracy on each corrupted test set and the associated ID clean accuracy (after OOD adaptation). The model performs lifelong adaptation, in which the model parameters will never be reset.
Figure G: Comparisons of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50. We record the OOD corruption accuracy on each corrupted test set and the associated ID clean accuracy (after OOD adaptation). The model performs single-domain adaptation, where model parameters of both Tent and our EATA are reset before adapting to a new corruption type.